### v1

@@ 5 Proof: existence and uniqueness of solutions
Let $(\rm H_{0})$} and $(\rm H_{q})$} holds and suppose $\xi$ is Malliavin differentiable with bounded derivative $\left\|D\xi\right\|_{\mathbb{S}^{\infty}}=\sup_{0\leq t\leq T}\left\|D_{t}\xi\right\|_{L^{\infty}}<\infty\text{ almost surely}$, then the mean-field BSDE admits a solution in $\mathbb{S}^{\infty}\times\mathbb{S}^{\infty}.$
@@
for all $k\in\mathbb{N}$ and $(t,z,z^{\prime},\tilde{z},\tilde{z}^{\prime})\in[0,T]\times(\mathbb{R}^{d})^{4}$. Thus, the corresponding BSDE with driver $\bar{g}_{k}$ satisfies $(\rm H_{0})$} and $(H_{\mathrm{lip}})$. Then, the mean-field BSDE
@@
We are now in a position to state the main result of this section. For the reader’s convenience, we restate the theorem below. To complete the proof of following theorem, it remains to extend the existence result from Malliavin differentiable terminal conditions to the general bounded case. Uniqueness then follows naturally from the stability result established in Theorem 4.6.
@@
To ensure that the approximating BSDEs satisfy Assumption $(\mathrm{H}_{0})$, we truncate each $\Phi^{n}$ by setting $\Phi^{n}:=\Phi^{n}\wedge\left\|\xi\right\|_{L^{\infty}}$. Hence, each $\xi^{n}$ remains bounded, and $(\xi^{n},g)$ satisfies $(\mathrm{H}_{0})$ for all $n\in\mathbb{N}$. By Lemma 5.1, each approximating BSDE with terminal condition $\xi^{n}$ admits a solution $(Y^{n},Z^{n})\in\mathbb{S}^{2}\times\mathbb{S}^{\infty}$, satisfying
@@
Furthermore, the stability result in Theorem 4.6 implies that (pick $p=4$), we have
$\|Y^{n}-Y^{m}\|_{\mathbb{S}^{2}}^{2}+\|Z^{n}-Z^{m}\|_{\mathbb{H}^{2}}^{2}\ \leq\ C_{\mathrm{global}}\;\|\xi^{1}-\xi^{2}\|_{L^{4}},$
where $C_{\mathrm{global}}$ depends only on $T$, $K_{z}$, $K_{\tilde{z}}$, and $K_{0}$, thus are independent of $n$ and $m$.
Since $(\xi^{n})_{n\in\mathbb{N}}$ is Cauchy in $L^{4}$, it follows that $(Y^{n},Z^{n})_{n\in\mathbb{N}}$ is also Cauchy in $\mathbb{S}^{2}\times\mathbb{H}^{2}$. As each $(Y^{n},Z^{n})$ lies in the Banach space $\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$, we conclude the existence of a limit $(Y,Z)\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$. Passing to the limit in (5.35), we verify that the pair $(Y,Z)$ indeed solves the original mean-field BSDE (2.5), establishing existence. Uniqueness then follows directly from the stability result in Theorem 4.6, which concludes the proof. ∎
@@ 6 Application: Utility Maximization
This section is structured as follows: In Section 6.1, we introduce the utility maximization problem and the associated mathematical framework. We then extend the given constraint set to incorporate a dependence on the stochastic process. It is important to note that, at this stage, we do not make any assumptions about this process beyond considering it as a general stochastic process. Next, we establish the connection between the utility maximization problem and the BSDE (6.37) with a thoughtfully chosen driver. We will subsequently demonstrate that the process defined earlier for the constraint set is precisely the solution $Z$ in the BSDE (6.37). With the complete framework in place, we offer a financial interpretation of the components, which are discussed in Remarks 6.4.
To connect this application to our earlier analysis about BSDE, we generalize the separately quadratic driver from BSDE (2.5) to fully coupled quadratic driver from BSDE (6.39) and establish the existence and uniqueness of solutions in Proposition 6.7. To do this we adopte a localization method in 6.6 then we extend it to the global case. Then this is the proof part for BSDE, then we connect this back to the utility maximization problem by presenting the final result in Theorem 6.9, we first derive a key estimate in Lemma 6.3. This estimate is then used in Theorem 6.9 to verify that the driver of BSDE (6.37) satisfies conditions $(\rm H_{0})$} and $(\rm H_{q}^{\prime})$}, ensuring the well-posedness of BSDE (6.39). Leveraging these lemmas, we conclude with the main result in Theorem 6.9, which is supported by the proof of Theorem 7 in hu2005utility.
@@ 6.1 Financial market model setup
Consider a financial market composed of $d$ risky stocks. In this market, there are countably infinite number of homogeneous investors who trades dynamically between the risk-free bond and the risky assets. An investor, a representative of homogeneous investors, uses the following model to maximize the utility. The price process of the $i$-th stock, denoted by $S^{i}$, evolves according to the following stochastic differential equation:
@@
where $W$ is $d$-dimensional Brownian motion, $b^{i}$ is a real-valued predictable, bounded stochastic process, representing the drift, and $\sigma^{i}$ is an $\mathbb{R}^{d}$-valued predictable bounded stochastic process, representing the volatility. The volatility matrix $\sigma_{t}=(\sigma_{t}^{1},\dots,\sigma_{t}^{d})^{\top}$ is assumed to have full rank $d$.
The market price of risk $\theta_{t}$, also referred to as the risk premium, is defined by:
@@
where $\theta_{t}$ is an $\mathbb{R}^{n}$-valued bounded predictable process.
@@
For the investors other than the representative, we assume the risk aversion parameter $\alpha>0$ are identical. The model for the risky assets are of the same type but with $b,\sigma,W$ being independent and identically distributed to the representative’s model. In addition, we assume the terminal liabilities are independent and has distribution identical to $F$.
@@ 6.2 Constraint set and aggregated market signal
To ensure that no arbitrage opportunities exist, the definition of admissible trading strategies imposes constraints on the set of allowable positions. Below, we outline the constraint set in Definition 6.1, which accounts for market-wide influences. To incorporate for the influence of market-wide behavior, we denote $\tilde{Z}$ for the trading strategy adjusted by $\theta/\alpha$, which is an $\mathbb{R}^{1\times d}$-valued stochastic process that satisfies the necessary integrability conditions. More precisely, $\tilde{Z}_{t}\sigma_{t}+\frac{\theta_{t}}{\alpha}$ yields the optimal strategy at time $t\in[0,T]$. In addition, for $t\in[0,T]$ and $\omega\in\Omega$, we define
$Z_{t}(\omega):=\tilde{Z_{t}}(\omega)\sigma_{t}(\omega),$
where $Z_{t}(\omega)$ is an element of $\mathbb{R}^{1\times n}$. These two stochastic processes are essentially equivalent because the matrix $\sigma$ has full rank. Later, we will show this $Z$ will precisely corresponds to the martingale density process of BSDE (6.37). We now present the following expressions for the constraint set under different settings. Specifically, we denote the constraint without market influence by $\tilde{D}\subseteq\mathbb{R}^{d}$, where $\tilde{D}$ is a closed set. When market influence depending on $\mathbb{E}[\tilde{Z}$] is present, the strategy $\pi_{t}(\omega)$ is constrained to $\tilde{C}(\mathbb{E}[\tilde{Z}_{t}])$, where $\tilde{C}(\mathbb{E}[\tilde{Z}_{t}])\subseteq\mathbb{R}^{d}$ is also a closed set.
The expectation of $\tilde{Z}$ is referred to as the aggregated market signal, representing the expected collective adjusted trading strategy of all market participants. This signal plays a crucial role in defining the feasible trading region (constraint set), as described below.
@@
Let $\tilde{D}$ be a closed set in $\mathbb{R}^{d}$, and define the mean-field adaptive constraint set as:
$\tilde{C}(\mathbb{E}[\tilde{Z}_{t}]):=\tilde{D}\odot L(\mathbb{E}[\tilde{Z}_{t}]),\quad t\in[0,T],$ (6.36)
where $\odot$ denotes the Hadamard product (component-wise multiplication) and $L(\mathbb{E}[\tilde{Z}_{t}])$ is a linear mapping depending on the value of $\mathbb{E}[\tilde{Z}_{t}]$ for fixed $t\in[0,T]$, representing how the trader adjusts decisions based on the market’s behavior, defined by:
$L(\mathbb{E}[\tilde{Z}_{t}]):=(\gamma\odot\mathbb{E}[\tilde{Z}_{t}])+\eta,$
@@
With the help of the previous definitions, for a fixed $t\in[0,T]$, we start with a set $\tilde{D}\in\mathbb{R}^{d}$, which represents certain constraints, such as risk limits or position constraints. Upon incorporating information from the market, we could apply scaling and shifting to this set, which now reflects both the original constraints and the market’s influence. This results in a new set that is much more flexible and realistic, capturing the dynamic behavior of the market. Therefore, we denote this updated set as $\tilde{C}(\mathbb{E}[\tilde{Z}_{t}])$, emphasizing its dependence on the aggregated market signal $\mathbb{E}[\tilde{Z_{t}}$] at time $t\in[0,T]$.
Definition 6.2 (Admissible strategies with constraints $\tilde{\mathcal{A}}$).
Let $\tilde{C}(\mathbb{E}[\tilde{Z}_{t}])$ be a closed set in $\mathbb{R}^{d}$ for fixed $t\in[0,T]$. The set of admissible trading strategies $\tilde{\mathcal{A}}$ consists of all $\mathbb{R}^{d}$-valued predictable processes $\pi$ satisfying $\pi_{t}\sigma_{t}\in L^{2}[0,T]$, $\pi_{t}\in\tilde{C}(\mathbb{E}[\tilde{Z}_{t}])$ a.s. $\forall t\in[0,T]$, and
@@
$V(0,x):=\sup_{\pi\in\tilde{\mathcal{A}}}\mathbb{E}\left[-\exp\left(-\alpha\left(x+\int_{0}^{T}\pi_{u}\frac{dS_{u}}{S_{u}}-F\right)\right)\right],$
where $V(0,\cdot)$ is called the value function at initial time 0. Solving this optimization problem requires selecting an admissible set from which we select the optimal trading strategy $\pi^{*}$.
To make this notation more compact, we introduce the following notations. For $t\in[0,T],\omega\in\Omega$, define the set $C_{t}\left(\omega,\mathbb{E}[Z_{t}]\right)\subseteq\mathbb{R}^{n}$ and (volatility-scaled) trading strategy $p_{t}$ such that
$C_{t}\left(\omega,\mathbb{E}[Z_{t}]\right):=\tilde{C}(\mathbb{E}[\tilde{Z}_{t}])\,\sigma_{t}(\omega)\quad\text{and}\quad p_{t}:=\pi_{t}\sigma_{t}.$
@@
Since the market sensitivity vector $\gamma$ and the market shift vector $\eta$ are both constant, the process $\sigma$ is uniformly bounded, and $\tilde{D}$ is closed, the result follows. ∎
Note that the admissible trading strategy set $\mathcal{A}$, which is equivalent to $\tilde{\mathcal{A}}$ (since $\sigma$ has full rank), consists of all $\mathbb{R}^{n}$-valued predictable processes $p$ that satisfy $p\in L^{2}[0,T]$ and $p_{t}(\omega)\in C_{t}\left(\omega,\mathbb{E}[Z_{t}]\right)$ almost surely for all $t\in[0,T]$. Additionally, the family
@@
is uniformly integrable. Finally, the dependency of the set $\tilde{C}(\mathbb{E}[\tilde{Z}_{t}])$ on the process $(\tilde{Z}_{t})_{0\leq t\leq T}$ is encoded by the dependency of the set $C_{t}\left(\omega,\mathbb{E}[Z_{t}]\right)$ on the process $(Z_{t})_{0\leq t\leq T}$. We keep the notation $X^{p}$ for the wealth process corresponding to investing strategy $p$.
@@
Earlier, we described the process $(Z_{t})_{0\leq t\leq T}$ simply as a stochastic process. However, in this context, we now define it as the martingale density process $(Z_{t})_{0\leq t\leq T}$ for the BSDE in equation (6.37). Consequently, as the constraint set depends on $\mathbb{E}[Z_{t}]$, the driver has dependence on $\mathbb{E}[Z_{t}]$: see (6.38).
Let us specify the driver $f$ for the construction of $\left\{V^{(p)}(\cdot,x):p\in\mathcal{A}\right\}$. By Itô’s formula,
@@
We have now completed the construction, and it is clear how the utility maximization problem connects to some types of BSDEs. In hu2005utility, the optimization problem was solved under the assumption that the random endowment $F$ is bounded and that the trading strategy $p$ takes values in the admissible set $\mathcal{A}$. However, this set differs from ours because, in their case, the constraint set for trading strategies is simply closed in $\mathbb{R}^{d}$. In contrast, we introduce a particular constraint set where the trader’s behavior depends on the market (i.e., the behavior of other traders) in a mean-field sense. Specifically, we expand the driver to include the expectation of the control process $Z$. Thus, we define a constraint set that depends on time, reflecting the market’s influence at each moment. This approach introduces greater flexibility and a more realistic model. To establish the connection between the control process $Z$ in the BSDE (6.37) and the trading strategy $p\in\mathcal{A}$, we provide the following remarks, which bridge the mathematical framework with financial interpretations.
@@
If we take $\tilde{D}$ to be the entire space, i.e., $\tilde{D}=\mathbb{R}^{d}$, then the resulting constraint set $\tilde{C}(\mathbb{E}[\tilde{Z}_{t}])$ is simply $\mathbb{R}^{d}$ as well. Consequently, we obtain $C_{t}\left(\omega,\mathbb{E}[Z_{t}]\right)=\mathbb{R}^{n}$, implying the absence of constraints. In this case, the optimal trading strategy is given by
@@
Thus, the process $Z_{t}=p_{t}^{*}-\frac{\theta_{t}}{\alpha}$ can be interpreted as the unconstrained trading strategy adjusted by $\frac{\theta_{t}}{\alpha}$ at time $t\in[0,T]$. In general, the constraint set $C$ depends on the average of the adjusted trading strategies among the market participants. The convergence is guaranteed by the homogeneity and exchangeability of market participants.

### v2

@@ 5 Proof: existence and uniqueness
Let $(\rm H_{0})$} and $(\rm H_{q})$} holds and suppose $\xi$ is Malliavin differentiable with bounded derivative $\left\|D\xi\right\|_{\mathbb{S}^{\infty}}<\infty$, then the mean-field BSDE admits a solution in $\mathbb{S}^{\infty}\times\mathbb{S}^{\infty}.$
@@
for all $k\in\mathbb{N}$ and $(t,z,z^{\prime},\tilde{z},\tilde{z}^{\prime})\in[0,T]\times[\mathbb{R}^{d}\times\mathbb{R}^{d}]^{2}$. Thus, the corresponding BSDE with driver $\bar{g}_{k}$ satisfies $(\rm H_{0})$} and $(H_{\mathrm{lip}})$. Then, the mean-field BSDE
@@
We are now in a position to state the main result of this section. For the reader’s convenience, we restate the main theorem below. To complete the proof of following theorem, it remains to extend the existence result from Malliavin differentiable terminal conditions to the general bounded case. Uniqueness then follows directly from the stability result established in Theorem 4.6.
@@
To ensure that the approximating BSDEs satisfy Assumption $(\mathrm{H}_{0})$, we truncate each $\Phi^{n}$ by setting $\Phi^{n}:=\Phi^{n}\wedge\left\|\xi\right\|_{L^{\infty}}$. Hence, each $\xi^{n}$ remains bounded, and $(\xi^{n},g)$ satisfies $(\mathrm{H}_{0})$ for all $n\in\mathbb{N}$. By Lemma 5.1, each approximating BSDE with terminal condition $\xi^{n}$ admits a solution $(Y^{n},Z^{n})\in\mathbb{S}^{\infty}\times\mathbb{S}^{\infty}$, satisfying
@@
Furthermore, the stability result in Theorem 4.6 implies that
$\|Y^{n}-Y^{m}\|_{\mathbb{S}^{2p}}+\|Z^{n}-Z^{m}\|_{\mathbb{H}^{p}}\ \leq\ C_{p}\;\|\xi^{n}-\xi^{m}\|_{L^{2p}}$
where $C_{p}$ depends only on $(T,K_{z},K_{\tilde{z}},K_{0})$, thus is independent of $n$ and $m$.
Pick $(\xi^{n})_{n\in\mathbb{N}}$ is Cauchy in $L^{2p}$, it follows that $(Y^{n},Z^{n})_{n\in\mathbb{N}}$ is also Cauchy in $\mathbb{S}^{2p}\times\mathbb{H}^{p}$. As each $(Y^{n},Z^{n})$ lies in the Banach space $\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$, we conclude the existence of a limit $(Y,Z)\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$. Passing to the limit in (5.28), we verify that the pair $(Y,Z)$ indeed solves the original mean-field BSDE (2.6), establishing existence. Uniqueness then follows directly from the stability result in Theorem 4.6, which concludes the proof. ∎
@@ 6 Application: Utility Maximization
This section is structured as follows: In Section 6.1, we introduce the utility maximization problem and the associated mathematical framework. We then extend the given constraint set to incorporate a dependence on the stochastic process. It is important to note that, at this stage, we do not make any assumptions about this process beyond considering it as a general stochastic process. Next, we establish the connection between the utility maximization problem and the BSDE (6.30) with a thoughtfully chosen driver. We will subsequently demonstrate that the process defined earlier for the constraint set is precisely the solution $Z$ in the BSDE (6.30). With the complete framework in place, we offer a financial interpretation of the components, which is discussed in Remarks 6.4.
To connect this application to our earlier analysis about BSDE, we generalize the separately quadratic driver from BSDE (2.6) to a fully coupled quadratic driver in BSDE (6.32) and establish the existence and uniqueness of solutions in Proposition 6.7 when the terminal condition is small enough. To do this, we adopt a localization method in 6.6, then we extend it to the global case. Then, after we established the existence and uniqueness for BSDE (6.32), we then connect this back to the utility maximization problem by presenting the final result in Theorem 6.9. We first derive a key estimate in Lemma 6.3. This estimate is then used in Theorem 6.9 to verify that the driver of BSDE (6.30) satisfies conditions $(\rm H_{0})$} and $(\rm H_{q}^{\prime})$}, ensuring the well-posedness of BSDE (6.32). Leveraging these lemmas, we conclude with the main result in Theorem 6.9, which is supported by the proof of Theorem 7 in hu2005utility.
@@ 6.1 Financial market model setup
Consider a financial market composed of $d$ risky stocks. In this market, there are countably infinite number of homogeneous investors who trade dynamically between the risk-free bond and the risky assets. An investor, a representative of homogeneous investors, uses the following model to maximize the utility. The price process of the $i$-th stock, denoted by $S^{i}$, evolves according to the following stochastic differential equation:
@@
where $W$ is $n$-dimensional Brownian motion, $b^{i}$ is a real-valued predictable, bounded stochastic process, representing the drift, and $\sigma^{i}$ is an $\mathbb{R}^{n}$-valued predictable bounded stochastic process, representing the volatility. The volatility matrix $\sigma_{t}=(\sigma_{t}^{1},\dots,\sigma_{t}^{d})^{\top}$ is assumed to have full rank $d$. And we assume that $\sigma_{t}\sigma_{t}^{\top}$ is uniformly elliptic. Thus market price of risk $\theta_{t}$, also referred to as the risk premium, is defined by:
@@
where $\theta_{t}$ is an $\mathbb{R}^{n}$-valued uniformly bounded predictable process.
@@
For the investors other than the representative, we assume the risk aversion parameter $\alpha>0$ is identical. The model for the risky assets is of the same type but with $b,\sigma,W$ being independent and identically distributed to the representative’s model. In addition, we assume the terminal liabilities are independent and have distribution $F$.
@@ 6.2 Constraint set and aggregated market signal
To ensure that no arbitrage opportunities exist, the definition of admissible trading strategies imposes constraints on the set of allowable positions. Below, we outline the constraint set in Definition 6.1, which accounts for market-wide influences. To incorporate the influence of market-wide behavior, for $t\in[0,T]$ and $\omega\in\Omega$, we define two stochastic processes that satisfy the necessary integrability conditions such that
$Z_{t}(\omega):=\hat{Z_{t}}(\omega)\sigma_{t}(\omega),$
where $Z_{t}(\omega)\in\mathbb{R}^{1\times n}$ and $\hat{Z_{t}}(\omega)\in\mathbb{R}^{1\times d}$. These two stochastic processes are essentially equivalent because the matrix $\sigma$ has full rank. Later on, in Remark 6.4, we will see $\hat{Z}_{t}\sigma_{t}+\frac{\theta_{t}}{\alpha}=Z_{t}+\frac{\theta_{t}}{\alpha}$ yields the optimal strategy at time $t\in[0,T]$ if no constants are imposed. Thus $Z$ can be interpreted as the trading strategy adjusted by $\theta/\alpha$. Later, we will show that this $Z$ will precisely correspond to the martingale density process of BSDE (6.30) as well.
We now present the following expressions for the constraint set under different settings. Specifically, we denote the constraint without market influence by $\hat{D}\subseteq\mathbb{R}^{d}$, where $\hat{D}$ is a closed set. When market influence is present, the strategy $\pi_{t}(\omega)$ is constrained to a new closed set.
The expectation of $\hat{Z}$ is referred to as the aggregated market signal, representing the expected collective adjusted trading strategy of all market participants. This signal plays a crucial role in defining the feasible trading region (constraint set), as described below.
@@
Let $\hat{D}$ be a closed set in $\mathbb{R}^{d}$, and define the mean-field adaptive constraint set as:
$\hat{C}(\mathbb{E}[\hat{Z}_{t}]):=\hat{D}\odot L(\mathbb{E}[\hat{Z}_{t}]),\quad t\in[0,T],$ (6.29)
where $\odot$ denotes the Hadamard product (component-wise multiplication) and $L(\mathbb{E}[\hat{Z}_{t}])$ is a linear mapping depending on the value of $\mathbb{E}[\hat{Z}_{t}]$ for fixed $t\in[0,T]$, representing how the trader adjusts decisions based on the market’s behavior, defined by:
$L(\mathbb{E}[\hat{Z}_{t}]):=(\gamma\odot\mathbb{E}[\hat{Z}_{t}])+\eta,$
@@
With the help of the previous definitions, for a fixed $t\in[0,T]$, we start with a set $\hat{D}\subset\mathbb{R}^{d}$, which represents certain constraints, such as risk limits or position constraints. Upon incorporating information from the market, we could apply scaling and shifting to this set, which now reflects both the original constraints and the market’s influence. This results in a new set that is much more flexible and realistic, capturing the dynamic behavior of the market. Therefore, we denote this updated set as $\hat{C}(\mathbb{E}[\hat{Z}_{t}])$, emphasizing its dependence on the aggregated market signal $\mathbb{E}[\hat{Z_{t}}$] at time $t\in[0,T]$.
Definition 6.2 (Admissible strategies with constraints $\hat{\mathcal{A}}$).
Let $\hat{C}(\mathbb{E}[\hat{Z}_{t}])$ be a closed set in $\mathbb{R}^{d}$ for fixed $t\in[0,T]$. The set of admissible trading strategies $\hat{\mathcal{A}}$ consists of all $\mathbb{R}^{d}$-valued predictable processes $\pi$ satisfying $\pi_{t}\sigma_{t}\in L^{2}[0,T]$, $\pi_{t}\in\hat{C}(\mathbb{E}[\hat{Z}_{t}])$ a.s. $\forall t\in[0,T]$, and
@@
$V(0,x):=\sup_{\pi\in\hat{\mathcal{A}}}\mathbb{E}\left[-\exp\left(-\alpha\left(x+\int_{0}^{T}\pi_{u}\frac{dS_{u}}{S_{u}}-F\right)\right)\right],$
where $V(0,\cdot)$ is called the value function at initial time 0. Solving this optimization problem requires selecting an admissible set from which we select the optimal trading strategy $\pi^{*}$. To make this notation more compact, we introduce the following notations. For $t\in[0,T],\omega\in\Omega$, define the set $C_{t}\left(\omega,\mathbb{E}[Z_{t}]\right)\subseteq\mathbb{R}^{n}$ and (volatility-scaled) trading strategy $p_{t}$ such that
$C_{t}\left(\omega,\mathbb{E}[Z_{t}]\right):=\hat{C}(\mathbb{E}[\hat{Z}_{t}])\,\sigma_{t}(\omega)\quad\text{and}\quad p_{t}:=\pi_{t}\sigma_{t}.$
Note $Z_{t}=\hat{Z_{t}}\sigma_{t}$ and $\mathbb{E}[Z_{t}]=\mathbb{E}[\hat{Z_{t}}]\sigma_{t}$.
@@
Since the market sensitivity vector $\gamma$ and the market shift vector $\eta$ are both constant, the process $\sigma$ is uniformly bounded, and $\hat{D}$ is closed, the result follows. ∎
Note that the admissible trading strategy set $\mathcal{A}$, which is equivalent to $\hat{\mathcal{A}}$ ($\sigma$ has full rank), consists of all $\mathbb{R}^{n}$-valued predictable processes $p$ that satisfy $p\in L^{2}[0,T]$ and $p_{t}(\omega)\in C_{t}\left(\omega,\mathbb{E}[Z_{t}]\right)$ almost surely for all $t\in[0,T]$. Additionally, the family
@@
is uniformly integrable. Finally, the dependency of the set $\hat{C}(\mathbb{E}[\hat{Z}_{t}])$ on the process $(\hat{Z}_{t})_{0\leq t\leq T}$ is encoded by the dependency of the set $C_{t}\left(\omega,\mathbb{E}[Z_{t}]\right)$ on the process $(Z_{t})_{0\leq t\leq T}$. We keep the notation $X^{p}$ for the wealth process corresponding to investing strategy $p$.
@@
Earlier, we described the process $(Z_{t})_{0\leq t\leq T}$ simply as a stochastic process. However, in this context, we now define it as the martingale density process $(Z_{t})_{0\leq t\leq T}$ for the BSDE in equation (6.30). Consequently, as the constraint set depends on $\mathbb{E}[Z_{t}]$, the driver should has dependence on $\mathbb{E}[Z_{t}]$ as well: see (6.31). To verify this, let us specify the driver $f$ for the construction of $\left\{V^{(p)}(\cdot,x):p\in\mathcal{A}\right\}$. By Itô’s formula,
@@
We have now completed the construction, and it is clear how the utility maximization problem in our setting connects to a type of mean-field type BSDEs as in hibon2017MFquadratic. In hu2005utility, the optimization problem was solved under the assumption that the random endowment $F$ is bounded and that the trading strategy $p$ takes values in the admissible set $\mathcal{A}$. However, this set differs from ours because, in their case, the constraint set for trading strategies is simply closed in $\mathbb{R}^{d}$. In contrast, we introduce a particular constraint set where the trader’s behavior depends on the market (i.e., the behavior of other traders) in a mean-field sense. Specifically, we expand the driver to include the expectation of the control process $Z$. Thus, we define a constraint set that depends on time, reflecting the market’s influence at each moment. This approach introduces greater flexibility and a more realistic model. To establish the connection between the control process $Z$ in the BSDE (6.30) and the trading strategy $p\in\mathcal{A}$, we provide the following remarks, which bridge the mathematical framework with financial interpretations.
@@
If we take $\hat{D}$ to be the entire space, i.e., $\hat{D}=\mathbb{R}^{d}$, then the resulting constraint set $\hat{C}(\mathbb{E}[\hat{Z}_{t}])$ is again $\mathbb{R}^{d}$ as well. Consequently, we obtain $C_{t}\left(\omega,\mathbb{E}[Z_{t}]\right)=\mathbb{R}^{n}$, implying the absence of constraints. In this case, the optimal trading strategy is given by
@@
Thus, the process $Z_{t}=p_{t}^{*}-\frac{\theta_{t}}{\alpha}$ can be interpreted as the unconstrained trading strategy adjusted by $\frac{\theta_{t}}{\alpha}$ at time $t\in[0,T]$. In general situation, the constraint set $C_{t}\left(\omega,\mathbb{E}[Z_{t}]\right)$ depends on the average of the adjusted trading strategies among the market participants. The convergence is guaranteed by the homogeneity and exchangeability of market participants.